### 16 Feb 2024 v3

# Current-induced near-field radiative energy, linear-momentum, and angular-momentum transfer

Abstract
In this paper, we study the near-field radiative energy, linear-momentum, and angular-momentum transfer from a current-biased graphene to nanoparticles. The electric current through the graphene sheet induces nonequilibrium fluctuations, causing energy and momentum transfer even in the absence of a temperature difference. The inherent spin-momentum locking of graphene surface plasmons leads to an in-plane torque perpendicular to the direction of the electric current. In the presence of a temperature difference, the energy transfer is greatly enhanced while the lateral force and torque remain within the same order. Our work explores the potential of utilizing current-biased graphene to manipulate nanoparticles.
I Introduction
The inclusion of nonreciprocal effects in the study of near-field thermal radiation has led to the discovery of various novel phenomena and has attracted significant scientific interest Moncada-Villa et al. (2015); Ben-Abdallah (2016); Zhu and Fan (2016); Latella and Ben-Abdallah (2017); Abraham Ekeroth et al. (2018); Ott and Biehs (2020); Tang et al. (2021a, b); Biehs et al. (2021). Nonreciprocity can be achieved in various scenarios. One example is by applying a magnetic field to magneto-optic materials Moncada-Villa et al. (2015); Ben-Abdallah (2016); Zhu and Fan (2016); Latella and Ben-Abdallah (2017); Abraham Ekeroth et al. (2018); Ott and Biehs (2020). Another approach is to use magnetic Weyl semimetals that have intrinsic time-reversal symmetry breaking Zhao et al. (2020); Guo et al. (2020); Tsurimaki et al. (2020); Pajovic et al. (2020).
Nonreciprocity-induced lateral force has been studied recently Shapiro (2017); Gelbwaser-Klimovsky et al. (2021); Khandekar et al. (2021); Tsurimaki et al. (2023). An atom or nanoparticle placed above a sample with mobile carrier drift experiences a fluctuation force Shapiro (2017). In a system where vacuum-separated plates are at different temperatures, if at least one plate is nonreciprocal material, the setup can function as a heat engine where thermal energy is converted into mechanical work Gelbwaser-Klimovsky et al. (2021). When a dipolar particle is placed in the vicinity of nonreciprocal plasmonic slab, it experiences a lateral force and torque in the presence of a temperature difference between the particle and the slab Khandekar et al. (2021). The force is linked to the nonreciprocity of the surface polaritons and the torque to the spin-momentum locking of the surface polaritons.
Current-biased graphene is also known for its nonreciprocity Borgnia et al. (2015); Duppen et al. (2016); Morgado and Silveirinha (2018, 2020). The exceptional electron mobility of graphene allows it to sustain nonreciprocal surface plasmon polaritons (SPPs) when driven by an electric current, which gives rise to interesting properties such as Fizeau drag Dong et al. (2021); Zhao et al. (2021) and negative Landau damping Morgado and Silveirinha (2017). Current-biased graphene is in a nonequilibrium state and this leads to a finite photonic chemical potential Henry and Kazarinov (1996) that depends on the in-plane wave vector for the thermal electromagnetic radiation. The occupation number of radiative photons is nonreciprocal, that is asymmetric with respect to the positive and negative in-plane wave vectors Volokitin and Persson (2008, 2011a, 2011b); Peng and Wang . The nonreciprocal properties of current-biased graphene can be adjusted through chemical doping or gate voltage, making it suitable for a wide range of applications in controlling thermal radiation. For example, near-field radiative heat transfer has been studied between two suspended graphene sheets Zhang et al. (2020a), graphene covered substrates Shi et al. (2022), graphene-nanoparticles Zhang et al. (2020b), and graphene-based multilayer systems Zhou et al. (2020). However, the physical aspects of the force and torque experienced by the nanoparticle near the current-biased graphene remain unexplored to date.
In this paper, we investigate the near-field transfer of energy, linear momentum, and angular momentum between an isotropic dipolar nanoparticle and the current-biased graphene. The nonequilibrium state in the current-biased graphene induces energy transfer between the nanoparticle and the graphene even in the absence of a temperature difference. Importantly, these SPPs carry both linear and spin-locked angular momentum. As a result, there is an exchange of net linear and spin angular momentum between the particle and the SPPs, leading to both lateral force and torque experienced by the nanoparticle.
Figure 1: (a) Schematic plot of the lateral force and torque exerted on an isotropic dipolar particle above a current-biased graphene in the near-field regime. The drift velocity $v_{d}$ is along the $x$ direction. (b) Dispersion of the SPPs of the current-biased graphene. The electric current effectively drags the SPPs which leads to the property of nonreciprocal propagation. (c) The photon occupation number difference $\delta n(\omega,q_{x})$ vs $\omega$ with $v_{d}=0.7v_{F}$.
II System and formalism
The system is illustrated schematically in Fig. 1(a), where the spherical nanoparticle with radius $R$ is placed near the surface of current-biased monolayer graphene. Without losing generality, we assume that the drift velocity $v_{d}$ in graphene is along the positive direction of the $x$ axis. Within the framework of fluctuational electrodynamics, the total power $H$, lateral force along the $x$ direction $F_{x}$, and torque along the $y$ direction $M_{y}$ can be expressed as
$P=\int_{0}^{\infty}\frac{d\omega}{2\pi}p(\omega),$ (1)
with $P\in\{H,F_{x},M_{y}\}$ and the corresponding spectral densities $p\in\{h,f_{x},m_{y}\}$. The spectral densities are given by Khandekar et al. (2021)
$h(\omega)$ $=-\hbar\omega k_{0}^{2}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (2)
$f_{x}(\omega)$ $=\hbar k_{0}^{2}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}q_{x}\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (3)
$m_{y}(\omega)$ $=-\hbar{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}q_{x}{\rm Im}\big{(}r_{p}e^{2i\beta_{0}d}\big{)}\delta n(\omega,q_{x}),$ (4)
with
$\Theta(\omega,{\bm{q}})={\rm Re}\bigg{[}\frac{r_{p}e^{2i\beta_{0}d}}{2\beta_{0}}\left(\frac{2q^{2}}{k_{0}^{2}}-1\right)\bigg{]},$ (5)
where the Fresnel reflection coefficient for $p$ polarization is given by
$r_{p}=\frac{\beta_{0}\sigma_{g}}{2\epsilon_{0}\omega+\beta_{0}\sigma_{g}}.$ (6)
Here, $\sigma_{g}$ is the surface conductivity of graphene, and $\epsilon_{0}$ is the dielectric constant in vacuum. The in-plane wave vector and the angular frequency are denoted by ${\bm{q}}=(q_{x},q_{y})$ and $\omega$. The magnitude of the out-of-plane wave vector in air is $\beta_{0}=\sqrt{k_{0}^{2}-q^{2}}$ with $k_{0}=\omega/c$ and $q=|{\bm{q}}|$. When an electric current is applied to the graphene so that the electrons near the Fermi surface experience a drift velocity, the system is not in local thermal equilibrium and the occupation number of the photons radiated from graphene is Doppler shifted. By assuming a drift velocity $v_{d}$ in the $x$ direction, the photon-occupation difference between the graphene and the nanoparticle is Volokitin and Persson (2008, 2011a, 2011b); Peng and Wang ; Tang et al. (2021b)
$\delta n(\omega,q_{x})=[e^{\hbar(\omega-q_{x}v_{d})/k_{B}T_{e}}-1]^{-1}-[e^{\hbar\omega/k_{B}T_{p}}-1]^{-1},$ (7)
where $T_{e}$ and $T_{p}$ are the temperatures of graphene and the nanoparticle, respectively. In getting the occupation number of the photons radiated from the current-biased graphene, one needs to apply the condition that $\hbar q_{x}v_{d}$ is much smaller than the chemical potential Peng and Wang . Equation (7) indicates that the electric current-induced nonequilibrium fluctuations can produce a finite energy flux even in the absence of a temperature difference. In the following, we assume that the temperature of the particle and the graphene is the same with $T_{p}=T_{e}=T=300\,$K, unless there is a special account for the temperatures. The energy transmission function of the energy flux in Eq. (2) is defined through $h(\omega)=\int d^{2}{\bm{q}}{\cal Z}_{h}(\omega,{\bm{q}})/4\pi^{2}$ with
${\cal Z}_{h}(\omega,{\bm{q}})=-\hbar\omega k_{0}^{2}{\rm Im}[\alpha(\omega)]\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (8)
which provides the information of the energy transfer at given $\omega$ and $\bm{q}$. Similar definitions are, respectively, applied to the force and torque.
[…]
The dispersion relation of SPPs in a monolayer graphene can be obtained through Gonçalves and Peres (2016)
$2\varepsilon_{0}\omega+\beta_{0}\sigma_{g}=0,$ (9)
where the nonlocal surface conductivity $\sigma_{g}$ is given by Svintsov and Ryzhii (2019)
$\sigma_{g}(\omega,q_{x},q_{y})$ $=\frac{ie_{0}^{2}\omega}{q^{2}}\frac{\mu(T)}{(\pi\hbar v_{F})^{2}}\int_{0}^{2\pi}d\theta\frac{1}{(1-\cos\theta v_{d}/v_{F})^{2}}\times$
$\frac{q_{x}(\cos\theta-v_{d}/v_{F})+q_{y}\sin\theta}{(\hbar\omega+i\gamma)/(\hbar v_{F})-q_{x}\cos\theta-q_{y}\sin\theta},$ (10)
with $\mu(T)=2k_{B}T\ln[2\cosh(\mu_{g}/2k_{B}T)]$. Here, $T$ is the temperature, $\mu_{g}$ is the chemical potential, $\gamma$ the damping parameter, and $e_{0}$ the electron charge. The nonlocal surface conductivity is obtained under the long-wavelength limit and by neglecting the interband transition which is justifiable for $2\mu_{g}>\hbar\omega$. In the numerical calculation, we set the Fermi velocity $v_{F}=10^{6}\,$m/s, the chemical potential $\mu_{g}=0.1\,$eV and the damping parameter $\gamma=3.7\,$meV. Figure 1(b) shows the nonreciprocal dispersions of SPPs in the presence of an electric current.
III Numerical Results
Before considering the nonequilibrium energy and momentum transfer, we discuss the nonreciprocal properties of $\delta n(\omega,q_{x})$ shown in Fig. 1(c). We can see that the photon occupation number difference $\delta n(\omega,q_{x})$ is positive for positive $q_{x}$ under $\omega{>}q_{x}v_{d}$. This indicates that the effective photonic temperature of the graphene is higher than the particle temperature $T_{p}$. Furthermore, it should be noted that contributions from $\omega{<}q_{x}v_{d}$ can be disregarded. This is because these contributions become significant only when the separation between the particle and the graphene reaches subnanometer scales Volokitin and Persson (2011b), whereas in our case, the distance is on the order of micrometers. Consequently, photons, especially those with positive $q_{x}$, flow from the graphene to the particle. In contrast, for negative $q_{x}$, $\delta n(\omega,q_{x})$ is always negative. Therefore, the effective photonic temperature of the graphene is lower than $T_{p}$. This sign reversal indicates a reversal in the direction of photon transfer compared to the behavior exhibited by positive $q_{x}$.
We consider three kinds of nanoparticle: NaCl, undoped GaAs, and InSb. The intricate dielectric characteristics are given by the Lorentz-Drude model with
$\varepsilon(\omega)=\varepsilon_{\infty}\frac{\omega_{L}^{2}-\omega^{2}-i\Gamma\omega}{\omega_{T}^{2}-\omega^{2}-i\Gamma\omega},$ (11)
where $\varepsilon_{\infty}$ is the high-frequency dielectric constant, and $\omega_{L}$ is the longitudinal optic-phonon frequency, $\omega_{T}$ is the transverse optic-phonon frequency, $\Gamma$ is the optic-phonon damping constant. Within the electrostatic limit, we can reasonably neglect multiple scattering interactions between the particle and the graphene, and describe the particle effectively using its electrical polarizability which is expressed as $\alpha(\omega)=4\pi R^{3}[\varepsilon(\omega)-1)]/[\varepsilon(\omega)+2]$ Zhao et al. (2012). It is important to note that these assumptions are valid only under the condition of $R\ll d\ll\hbar c/k_{B}T$. In Fig. 2(a), we present the imaginary part of the polarizability for these nanoparticles, revealing distinct dipolar resonances.
Now we study the effect of drift velocity on the net power transfer experienced by different nanoparticles, as shown in Fig. 2(b). We differentiate between positive and negative energy fluxes, which represent the flow of energy from the graphene to the particle and from the particle to the graphene, respectively. Our analysis reveals intriguing behaviors. For NaCl and InSb particles, the power transfer initially increases as $v_{d}$ increases. However, with increasing $v_{d}$ further, the power transfer decreases. In contrast, the direction of net power transfer undergoes a change for GaAs particles as the drift velocity increases. These are due to the interplay between the nonreciprocity of the SPPs and photon occupation number of graphene.
[…]
To analyze the contributions of forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs to net power transfer, we use the NaCl particle as an example. In Fig. 2(c), we show the energy transmission function and the dispersion of graphene plasmons (shown by the cyan dashed line). This visual representation aligns the bright branches of the energy transmission function with the cyan dashed line, confirming the dominant role played by SPPs in governing power transfer. For $q_{x}{>}0$, the energy transfer coefficients are positive, indicating energy transfer from graphene to the particle. For $q_{x}{<}0$, the energy transfer coefficients are negative, signifying energy transfer from the particle back to the graphene. The net energy transfer is the sum of contributions from $q_{x}{>}0$ and $q_{x}{<}0$. At lower drift velocities, the nonreciprocal characteristics of SPPs are not pronounced, and energy transfer is mainly governed by the nonreciprocal photon occupation number difference $\delta n(\omega,q_{x})$, as shown in Fig. 1(c). However, as the drift velocity increases, nonreciprocal SPPs interplay with $\delta n(\omega,q_{x})$. The high drift velocities suppress the net power transfer due to the nonreciprocal photon occupation number even change the energy transfer direction between GaAs and graphene.
In the following, we investigate the transfer of linear and angular momenta between the nanoparticles and the current-biased graphene. We define positive $M_{y}$ along the negative $y$ direction and positive $F_{x}$ as oriented in the positive $x$ direction. Our findings revealed a nonmonotonic behavior: Both force [Fig. 3(a)] and torque [Fig. 3(b)] increase with increasing the drift velocity up to a certain value, then start to decrease. This behavior is consistent with the energy transfer.
Figure 4: (a) Net power and (b) net force for the NaCl particle as a function of $v_{d}$ at various temperatures $T_{e}$ in graphene with $T_{p}=300\,$K and $d=0.2\,\mu$m. The curve of power transfer at $T_{e}=300\,$K is multiplied by $10^{3}$ for better visibility. (c) Net power and (d) net force vs $v_{d}$ with $T_{e}=320\,$K and $T_{p}=300\,$K. The contributions from the forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs are separately shown.
To have a deeper understanding of the force and torque acting on the particle, we analyze the exchange of linear momentum and spin angular momentum between the forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs and the nanoparticle. When photons transfer from the graphene to the particle ($q_{x}{>}0$), the particle gains momentum, contributing to the lateral force along the positive $x$ direction. Conversely, when $q_{x}{<}0$, the particle loses momentum due to the reverse photon transfer, producing a lateral force in the same direction. Therefore, SPPs with both positive and negative $q_{x}$ generate a net lateral force directed along the positive $x$ direction. Due to the spin-momentum locking of SPPs Bliokh et al. (2015), the transverse spin of SPPs strongly couples with their propagation direction. As a result, when SPPs propagate in the positive (negative) $x$ direction, their spin aligns with the negative (positive) $y$ direction, respectively. This spin angular momentum transfer is facilitated by photon transfer from the graphene to the particle in the case of $q_{x}{>}0$, generating a torque directed along the negative $y$ direction. Importantly, the torque contributed from $q_{x}{<}0$ is also along the negative $y$ direction. Consequently, the net lateral torque is along the negative $y$ direction. These are supported by Fig. 3(c) which shows the torque spectral density $m_{y}$ as a function of $q_{x}$ and $q_{y}$ at $\hbar\omega=27.8\,$meV and $v_{d}=0.3v_{F}$. Notably, for both positive and negative $q_{x}$, the torque spectral densities $m_{y}$ are positive. Remarkably, the force and torque induced by current-biased graphene are of the same order of magnitude as those generated by thermally excited nonreciprocal surface electromagnetic waves Khandekar et al. (2021).
Finally, we consider the scenario where the temperature of the particle differs from that of the graphene. Figures 4(a) and  4(b) show the dependence of net energy transfer and lateral force on the NaCl particle at various temperatures of the graphene. The nanoparticle temperature is at 300300300\,300K. Notably, the energy transfer from the graphene to the nanoparticle increases by three orders of magnitude. To explain this, we calculate the contributions from both the forward and backward SPPs at $T_{e}{>}T_{p}$ [see Figs. 4(c) and (d)]. It can be seen from Fig. 4(c) that the energy transfer for photons with opposite momentum has the same sign, which is completely different from the case in the absence of a temperature difference. This means that the contribution from the temperature difference dominates over that from the electric current. Therefore, the energy transfer is greatly enhanced in the presence of a temperature difference. In contrast, the photons with $q_{x}{>}0$ and $q_{x}{<}0$ make positive and negative contributions to the force, respectively. We find that the force experienced by the particle does not differ significantly from that without a temperature difference. The torque has similar behavior as the force. Thus, the temperature difference has a marginal impact on the lateral force and torque while significantly enhancing the net power transfer with $T_{e}{>}T_{p}$.
IV Conclusion
We have investigated the near-field radiative energy, linear-momentum, and angular-momentum transfer from a current-biased graphene sheet to nanoparticles. Notably, for the GaAs particle, both the magnitude and the direction of energy transfer can be electrically manipulated. This unique phenomenon results from the interplay of nonreciprocal photon occupation differences and graphene surface plasmon polaritons. Moreover, in the presence of a temperature difference, our study reveals distinct energy transfer behavior while maintaining consistent force characteristics compared to the scenario without a temperature difference. Recently, measuring energy transfer, force, and torque between nanoparticles and surfaces has been experimentally realized Menges et al. (2016); Ju et al. (2023); Ahn et al. (2020), so our theoretical predictions can thus be realized in the near future.
Acknowledgements.
H.Z., L.Z., and J.C. acknowledge the support from the National Natural Science Foundation of China (Grants No. 12074230, No. 12174231), the Fund for Shanxi “1331 Project,” Fundamental Research Program of Shanxi Province through 202103021222001, Program of Education and Teaching Reform in Shanxi Province (Grant No. J20230003), and Research Project Supported by Shanxi Scholarship Council of China. G.T. is supported by National Natural Science Foun dation of China (Grants No. 12374048 and No. 12088101) and NSAF (Grant No. U2330401). This research was partially conducted using the High Performance Computer of Shanxi University.
$*{}^{*}$gmtang@gscaep.ac.cn
${}^{\dagger}$zhanglei@sxu.edu.cn
${}^{{\ddagger}}$chenjun@sxu.edu.cn
Appendix A Derivation of the spectral densities of the power transfer, the force and the torque
From the fluctuational electrodynamics, the spectral densities of the net power transfer from the particle to the environment $h$, the force $f_{j}$, and the torque $m_{j}$ can be expressed as Henkel et al. (2002); Manjavacas et al. (2017)
$h(\omega)$ $=\omega{\rm Im}\Big{[}\Big{\langle}p_{j}^{{\text{fl}}^{*}}(\omega)E_{j}^{\text{ind}}(\omega)+p_{j}^{{\text{ind}}^{*}}(\omega)E_{j}^{{\text{fl}}}(\omega)\Big{\rangle}\Big{]},$ (12)
$f_{j}(\omega)$ $={\rm Re}\Big{[}\Big{\langle}p_{k}^{\text{fl}}(\omega)\partial_{j}E_{k}^{\text{ind}^{*}}(\omega)+p_{k}^{\text{ind}}(\omega)\partial_{j}E_{k}^{{\text{fl}}^{*}}(\omega)\Big{\rangle}\Big{]},$ (13)
$m_{j}(\omega)$ $=\epsilon_{jkl}{\rm Re}\Big{[}\Big{\langle}p_{k}^{{\text{fl}}}(\omega)E_{l}^{{\text{ind}}^{*}}(\omega)+p_{k}^{\text{ind}}(\omega)E_{l}^{{\text{fl}}^{*}}(\omega)\Big{\rangle}\Big{]},$ (14)
[…]
The dipole moment induced by the electric-field fluctuations is given by
$p_{j}^{\text{ind}}(\bm{r})=\epsilon_{0}\alpha_{jk}E_{k}^{\mathrm{fl}}(\bm{r}),$ (15)
where $\alpha$ is the polarizability of the particle. The induced field by the dipole fluctuations can be expressed as
$E_{j}^{\mathrm{ind}}(\mathbf{r}_{2},\omega)=\omega^{2}\mu_{0}G_{jk}(\mathbf{r}_{2},\mathbf{r},\omega)p_{k}^{\mathrm{fl}}(\mathbf{r},\omega),$ (16)
where the Green’s function in the dipole-plate geometry is expressed as Khandekar and Jacob (2019)
$G(\mathbf{r}_{2},\mathbf{r}_{1},\omega)=\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}e^{i{\bm{q}}\cdot\big{(}\mathbf{R}_{2}-\mathbf{R}_{1}\big{)}}G(\mathbf{q},z_{1}+z_{2},\omega),$ (17)
with
$G(\mathbf{q},z_{1}+z_{2},\omega)=e^{i\beta_{0}(z_{1}+z_{2})}\frac{ir_{p}}{2\beta_{0}}\hat{\mathbf{e}}_{p+}\hat{\mathbf{e}}_{p-}^{T}.$ (18)
Here, we use the notation $\mathbf{r}_{n}=(\mathbf{R}_{n},z_{n})$ with $n=1,2$. The polarization vectors $\hat{\mathbf{e}}^{T}_{p\pm}$ for the $p$ polarized mode along the ±$\pm z$ direction are
$\hat{\mathbf{e}}^{T}_{p\pm}=\frac{1}{k_{0}}\left(\mp\beta_{0}\cos\phi,\ \mp\beta_{0}\sin\phi,\ q\right).$ (19)
The in-plane wave vector is ${\bm{q}}=(q_{x},q_{y})=q(\cos\phi,\sin\phi)$ where $\phi$ is the angle with respect to the $x$-axis. The out-of-plane wave vector in air is $\beta_{0}=\sqrt{k_{0}^{2}-q^{2}}$ with $k_{0}=\omega/c$. The Fresnel reflection coefficient for $p$ polarization is denoted as $r_{p}$.
The correlations functions of the fluctuating quantities, $p^{\mathrm{fl}}$ and $E^{\mathrm{fl}}$, respectively, satisfy the following fluctuation-dissipation relations with Landau and Lifshitz (2013)
$\Big{\langle}p_{j}^{\mathrm{fl}}(\mathbf{r},\omega)p_{k}^{\mathrm{fl}^{*}}(\mathbf{r},\omega^{\prime})\Big{\rangle}=$ $(1/2i)\hbar\epsilon_{0}\big{[}\alpha_{jk}(\omega)-\alpha_{kj}^{*}(\omega)\big{]}\times$
$\bigg{[}n(T_{p},\omega)+\frac{1}{2}\bigg{]}2\pi\delta(\omega-\omega^{\prime}),$ (20)
and
$\Big{\langle}E_{j}^{\mathrm{fl}}(\mathbf{r}_{1},\omega)E_{k}^{\mathrm{fl}^{*}}(\mathbf{r}_{2},\omega^{\prime})\Big{\rangle}=(1/2i)\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}\bigg{[}e^{i{\bm{q}}\cdot\big{(}\mathbf{R}_{1}-\mathbf{R}_{2}\big{)}}$
$G_{jk}(\mathbf{q},z_{1}+z_{2},\omega)-e^{i{\bm{q}}\cdot\big{(}\mathbf{R}_{2}-\mathbf{R}_{1}\big{)}}G_{kj}^{*}(\mathbf{q},z_{1}+z_{2},\omega)\bigg{]}\times$
$\Big{(}\hbar\mu_{0}\omega^{2}\Big{)}\bigg{[}n(T_{e},\omega-q_{x}v_{d})+\frac{1}{2}\bigg{]}2\pi\delta(\omega-\omega^{\prime})\delta(\mathbf{r}_{2}-\mathbf{r}_{1}),$ (21)
where $n(T,\omega)=\left[\exp(\hbar\omega/{k_{B}T})-1\right]^{-1}$ is the Bose-Einstein distribution function at temperature $T$. Here, the correlation functions are antisymmetrized. The temperatures of the nanoparticle and the graphene are, respectively, denoted as $T_{p}$ and $T_{e}$. The photon occupation number radiated from the graphene is Doppler shifted due to the electric current induced drift velocity $v_{d}$ in the $x$ direction Volokitin and Persson (2008, 2011a, 2011b); Peng and Wang ; Tang et al. (2021b). To get equal frequency correlations, the factor of $2\pi$ in front of the delta functions in the above correlation functions is canceled by integrating over frequency $\omega^{\prime}$.
The radiation power spectral density can be obtained as
$h(\omega)$ $=\omega{\rm Im}\Big{[}\omega^{2}\mu_{0}\Big{\langle}p_{j}^{\text{fl}^{*}}(\mathbf{r},\omega)p_{k}^{\text{fl}}(\mathbf{r},\omega)\Big{\rangle}G_{jk}(\mathbf{r},\mathbf{r},\omega)$
$\quad\quad\quad+\epsilon_{0}\alpha^{*}(\omega)\Big{\langle}E_{j}^{\text{fl}^{*}}(\mathbf{r},\omega)E_{j}^{\text{fl}}(\mathbf{r},\omega)\Big{\rangle}\Big{]}$
$=$ $-\frac{\hbar\omega^{3}}{c^{2}}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}{\rm Im}\big{[}G_{jj}(\mathbf{q},2d,\omega)\big{]}\delta n(\omega,q_{x}),$ (22)
with the photon-occupation difference
$\delta n(\omega,q_{x})=[e^{\hbar(\omega-q_{x}v_{d})/k_{B}T_{e}}-1]^{-1}-[e^{\hbar\omega/k_{B}T_{p}}-1]^{-1}.$ (23)
We define the energy transmission function ${\cal Z}_{h}(\omega,{\bm{q}})$ through
$h(\omega)=\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}{\cal Z}_{h}(\omega,{\bm{q}}),$ (24)
so that
${\cal Z}_{h}\left(\omega,{\bm{q}}\right)=-\hbar\omega k_{0}^{2}{\rm Im}\big{[}\alpha(\omega)\big{]}\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (25)
with
$\Theta(\omega,{\bm{q}})={\rm Re}\Big{[}r_{p}e^{2i\beta_{0}d}\big{(}2q^{2}-k_{0}^{2}\big{)}/(2\beta_{0}k_{0}^{2})\Big{]}.$ (26)
The spectral density of force can be obtained as
$*\displaystyle f_{x}(\omega)={\rm Re}\Big{[}\Big{\langle}p_{k}^{\text{fl}}(\mathbf{r},\omega)\partial_{x}\big{(}\omega^{2}\mu_{0}G_{kj}(\mathbf{r},\mathbf{r},\omega)p_{j}^{\text{fl}}(\mathbf{r},\omega)\big{)}^{*}$
$\quad\quad+\epsilon_{0}\alpha(\omega)E_{k}^{\text{fl}}\left(\mathbf{r},\omega\right)\partial_{x}E_{k}^{\text{fl}^{*}}(\mathbf{r},\omega)\Big{\rangle}\Big{]}$
$=$ $\hbar k_{0}^{2}{\rm Im}\big{[}\alpha(\omega)\big{]}\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}q_{x}{\rm Im}\big{[}G_{kk}(\mathbf{q},2d,\omega)\big{]}\delta n(\omega,q_{x}).$ (27)
In the above, $\partial_{x}$ represents the partial derivative in the $x$ direction of position $\mathbf{r}$. The spectral density of torque can be obtained as
$m_{y}(\omega)=$ $\omega^{2}\mu_{0}{\rm Re}\Big{[}\Big{\langle}p_{z}^{\text{fl}}(\mathbf{r},\omega)G_{xz}^{*}(\mathbf{r},\mathbf{r},\omega)p_{z}^{\text{fl}^{*}}(\mathbf{r},\omega)\Big{\rangle}$
$-\Big{\langle}p_{x}^{\text{fl}}(\mathbf{r},\omega)G_{zx}^{*}(\mathbf{r},\mathbf{r},\omega)P_{x}^{\text{fl}^{*}}(\mathbf{r},\omega)\Big{\rangle}\Big{]}$
$+\epsilon_{0}{\rm Re}\Big{[}\alpha(\omega)\Big{\langle}E_{z}^{\text{fl}}(\mathbf{r},\omega)E_{x}^{\text{fl}^{*}}(\mathbf{r},\omega)\Big{\rangle}$
$-\alpha(\omega)\Big{\langle}E_{x}^{\text{fl}}(\mathbf{r},\omega)E_{z}^{\text{fl}^{*}}(\mathbf{r},\omega)\Big{\rangle}\Big{]}$
$=$ $\hbar k_{0}^{2}{\rm Im}\big{[}\alpha(\omega)\big{]}\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}{\rm Re}\Big{[}G_{zx}(\mathbf{q},2d,\omega)$
$-G_{xz}(\mathbf{q},2d,\omega)\Big{]}\delta n(\omega,q_{x}).$ (28)
We thus have
$f_{x}(\omega)=\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}{\cal Z}_{f}(\omega,{\bm{q}}),\ m_{y}(\omega)=\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}{\cal Z}_{m}(\omega,{\bm{q}}),$ (29)
with the force and the torque transmission functions, respectively, as
${\cal Z}_{f}(\omega,{\bm{q}})=\hbar k_{0}^{2}q_{x}{\rm Im}\big{[}\alpha(\omega)\big{]}\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (30)
${\cal Z}_{m}(\omega,{\bm{q}})=-\hbar q_{x}{\rm Im}\big{[}\alpha(\omega)\big{]}{\rm Im}\big{(}r_{p}e^{2i\beta_{0}d}\big{)}\delta n(\omega,q_{x}).$ (31)